Barry spent 1/3 of his monthly salary for rent and 1/5 of his monthly salary for his car payment. If $784 was left, what was his monthly salary?

Answer by Boreal(15235)

He spent (1/3)+(1/5) total with a common denominator of 15, or (5/15)+(3/15)=(8/15)
(7/15) of his salary was left
(7/15)x=$784, if x is the salary.
784*(15/7)=$1680 is his monthly salary ANSWER
(1/3) spent was $560
(1/5 spent was $336
total is $896, which subtracted from $1680=$784.